Circular No. 8474

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005V IN NGC 2146 Further to IAUC 8473, S. Mattila et al. report the discovery of another

apparent supernova (mag 13.8) in K_{s}-band images obtained on Jan. 30.16 UT at the William Herschel Telescope. The new object is located at α = 6^{h}18^{m}38^{s}.28, δ = +78^{o}21′28′′.8 (equinox 2000.0), which is 1′′.8 east and 3′′.4 north of the K_{s}-band nucleus of the starburst galaxy NGC 2146. J- and H-band observations of SN 2005V yield preliminary colors J − H = +0.13 ± 0.33 and H − K = +0.18 ± 0.34, indicating modest extinction.

Following requests by the Central Bureau, T. Boles (Coddenham, Suffolk, England; limiting mag 20.5), T. Puckett (Ellijay, GA; limiting mag 20), and K. Kadota (Ageo, Saitama-ken, Japan; via S. Nakano) report that their CCD images obtained on Feb. 1 UT show that SN 2005V could not be definitively separated from the bright core of NGC 2146.

S. Taubenberger and A. Pastorello, Max-Planck-Institut für Astrophysik, Garching; and S. Benetti, Istituto Nazionale di Astrofisica, Osservatorio Astronomico di Padova, on behalf of the European RTN collaboration, report that inspection of a spectrogram of SN 2005V, obtained on Feb. 1.0 UT by J. Aceituno with the Calar Alto 2.2-m telescope (+ CAFOS; range 330–880 nm), shows it to be a type-Ib/c supernova, ∼ 1–2 weeks past maximum. The very red continuum and the deep interstellar Na I D (EW = 0.55 nm) suggest heavy host-galaxy extinction. The spectrum shows P-Cyg lines of Fe II, Na I D, O I, and Ca II. Also, a strong absorption at 620 nm is visible, possibly due to Si II. No clear evidence of He I lines is found.

SUPERNOVA 2005M IN NGC 2930 R. C. Thomas, Lawrence Berkeley National Laboratory, writes on be-

half of the ‘Nearby Supernova Factory’ that a spectrogram (range 330–990 nm) of SN 2005M (IAUC 8470), obtained Jan. 24.52 UT with the University of Hawaii 2.2-m telescope (+ SuperNova Integral Field Spectrograph), shows it to be a type-Ia supernova that is most similar to SN 1991T at ∼ 12 days before maximum. Weak Ca II H and K and Fe III features are evident. The spectrum is very blue and lacks interstellar Na I D, indicating very little or no extinction.

COMET 164P/2004 Y1 (CHRISTENSEN) Following the identification of observations in 1998 Jan. and Apr.

(cf. MPC 53463), comet P/2004 Y1 (cf. IAUC 8458) has been numbered 164P/Christensen.

2005 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green